### v1

# More About the Spontaneous Breaking of Time Reversal in de Sitter Space

More About the Spontaneous Breaking of Time Reversal in de Sitter Space
Leonard Susskind LITP and Department of Physics, Stanford University, Stanford, CA 94305-4060, USA Google, Mountain View, CA, USA
It is widely thought that the quantum theory of de Sitter space requires the existence of a physical observer in the static patch. What exactly is meant by an observer is unclear; it could be anything from a few photons with energy just above the Gibbons-Hawking temperature to a gravitationally bound cluster of galaxies. In a recent note I explained that the need for observers can arise from the spontaneous breaking of time-reversal symmetry. This longer paper expands on the subject, filling in conceptual arguments that were implicit but not explicitly stated in the earlier paper.
1 Introduction
There is a widespread belief that the quantum mechanics of de Sitter space only makes sense if there is an observer present [1]. Discrepancies between semiclassical gravity and static patch holography have been resolved by introducing a feature localized at the pode of the static patch. According to the terminology of [1][2][3][4] this feature is called an “observer.” I put “observer” in quotes because I don’t believe it necessarily has much to do with observation by almost classical apparatuses. “Observers” comprise a more general class than observers. Observers, among other requirements, are heavy enough to be localized near the pode; observers can be much lighter and can have wavelengths of cosmic scale.
The ability of an observer to resolve the discrepancies is surprising because the semiclassical calculation may involve bulk degrees of freedom localized extremely far from the pode—for example near the horizon—and yet no matter how far away the pode is from the region of the semiclassical calculation, the presence of an observer at the pode changes things enough to resolve the discrepancy. On the face of it this is a form of non-locality forbidden in bulk quantum field theory.
One thing about “observers” is that they should have some form of time-reversal asymmetry to distinguish future and past. I’ll call such an asymmetric feature a “clock.”
Generally the literature on this subject does not spell out exactly what is meant by an observer. In fact it seems that it could be anything from a galaxy to a tiny out-of-equilibrium feature in the Hawking radiation—no more than a handful of photons with energy of order the Gibbons-Hawking temperature. Calling such a feature an observer seems misleading to me; I prefer the terms feature, lump, or fluctuation to the overly suggestive observer but to avoid the introduction of new terminology I’ll stick with “observer.” But it is important to keep in mind that “observer” is just a name for an feature that could be almost anything.
To make matters more extreme, not only can the observer be extremely far from the region of the semiclassical calculation; it may also be a very feeble perturbation and still resolve the mismatch. It can be any “observer” as long as it can be localized in the bulk, far from the horizon. This means that its wavelength should be a small fraction of the de Sitter length $\ell.$
Now that I’ve explained the “observer” let’s turn to the region $h$ in figure 1. I’ll refer to it as the laboratory (without quotes). It is the region of interest where experiments take place, the experiments being described by a set of correlation functions localized near $h.$
In a recent note [5] I argued that the need for an “observer” has nothing to do with the consistency of quantum mechanics; it is due to a spontaneous symmetry breaking (SSB) of time-reversal symmetry. The “observer” is just the small feature which kicks off an instability of the symmetric vacuum and restores cluster decomposition [^1] Violations of cluster decomposition, aka infinite range correlations, are generally associated with spontaneous symmetry breaking..
In this followup the arguments of [5] will be clarified and made more complete.
I will use the notation $\cal{T}$} for time-reversal, both the symmetry and the antiunitary time-reversal operator. It has the following properties:
$\cal{T}$} is antiunitary. It is the product of a unitary operator and complex conjugation.
$\cal{T}$} satisfies
${\cal{T}}^{-1}={\cal{T}}.$
I’ll begin by reviewing the conclusions of [5]:
  1. 1.
    There is a breakdown of bulk cluster decomposition in the usual de Sitter vacuum. In the bulk theory the de Sitter vacuum means the Hartle-Hawking state; in the holographic boundary theory it is the maximally entangled thermofield-double state.
  2. 2.
    A failure to cluster is symptomatic of spontaneous symmetry breaking. For the case at hand the spontaneously broken symmetry is $\cal{T}$} (time-reversal).
  3. 3.
    When SSB takes place the symmetric vacuum is unstable to very slight perturbations which can kick the vacuum into one of the stable asymmetric states.
  4. 4.
    In the unstable $\cal{T}$}-symmetric vacuum correlation functions do not match semiclassical expectations; for example two-point functions are necessarily real in the symmetric vacuum but are complex in the semiclassical theory. In the stable asymmetric vacuum correlation functions are consistent with semiclassical expectations.
  5. 5.
    I will add one more thing to the list which may or may not have been clear from [5]. There is a subtlety about the meaning of time-reversal in de Sitter space; in ordinary quantum mechanics clocks may be thought of as registering absolute Newtonian time; for that reason we never need to introduce them explicitly. They live outside the system of interest and do not interact with it [^2] If the clock interacted with the system the system would back-react on the clock and the clock could not register absolute time..
    In a closed universe without boundaries, e.g., de Sitter space, the only physically meaningful time is that registered by a real material clock. This raises a question about the meaning of $\cal{T}$}:
    When we reverse time do we only reverse the motion of the system, or do we also reverse the clock?
    In ordinary QM we reverse only the system but not the clock. In de Sitter space we must also reverse the clock. I will explain the subtlety by example in what follows.
    Let me depart from the technical argument here and give an analogy. Imagine a quantum movie of a quantum oscillator and a real clock in the movie theater keeping time. Now run the movie backward. Is there really a symmetry or do we also have to run the clock backward? In principle these are two different operations. If the clock and the system are totally non-interacting there is no important difference between the two but if there are any interactions then only the combined time reversal is a symmetry.
Harlow and collaborators have argued that in quantum gravity $\cal{T}$} is a gauge symmetry [6]. I’ll discuss this issue, but in short I think it is important to consider it so.
In the next two sections I will illustrate some basic concepts about $\cal{T}$} and SSB using elementary examples. After that I will put them together in the context of de Sitter space.
2 Time-Reversal With or Without Clocks
Let’s begin with a system and a clock. I’ll choose a very simple system: the Harmonic oscillator and later a very simple clock.
2.1 The System
The equations are too well known to require explanation, I’ll just write them down for the case $\omega=1.$
$[a^{-},a^{+}]$ $=$ $1$ (2.1)
$H_{s}$ $=$ $a^{+}a^{-}$ (2.3)
$x(t)$ $=$ $\frac{a^{+}e^{it}++a^{-}e^{-it}}{2}$ (2.5)
The subscript $s$ means system.
By the usual definition of time-reversal the harmonic oscillator is a $\cal{T}$}-invariant system. Let’s test that out. The operator $C$
$C$ $=$ $[x(-s),x(s)]$ (2.6)
$x(s)$ $=$ $e^{iHs}xe^{-iHs}$ (2.8)
(for any fixed [^3] A more general version of (4.23) is C(s,t)=[x(ts),x(t+s)]C(s,t)=[x(t-s),x(t+s)] numerical value of $s$) changes sign under $\cal{T}$}; in other words it is $\cal{T}$}-odd. To see that we use
${\cal{T}}x(t){\cal{T}}=x(-t)$
and we find ${\cal{T}}C{\cal{T}}=-C.$ Now consider the expectation value of $C$ in any $\cal{T}$}-invariant state, for example the oscillator ground state. Since $C$ is odd and the state is even under $\cal{T}$} we might expect
$\langle 0|C|0\rangle=0.$
But explicit calculation gives,
$\langle 0|C|0\rangle=\frac{i}{2}\sin{s}$ (2.9)
Only the real part of $\langle 0|C|0\rangle$ vanishes for any $s\neq 0.$ Evidently the logic which says that $\langle 0|C|0\rangle$ vanishes is wrong.
2.2 The Clock
In ordinary quantum mechanics time is “Newtonian,” that is to say time recorded by an ideal clock that lives outside the system without interacting with it. We will denote it by $t.$ In the real world, especially in cosmology, any clocks that exist are part of the universe, the universe being the system we are studying. This is not important in ordinary quantum mechanics but in a closed universe like de Sitter space it is all important.
Let us introduce a simple model for a clock, namely a free particle with coordinate $\tau$ and conjugate momentum $\pi.$
$\tau,\pi]=i$ (2.10)
The clock Hamiltonian is,
$H_{c}=|\pi|$ (2.11)
The $\tau$ axis is the “face of the clock” and the $\tau$ operator represents clock time. Ideal Newtonian time is still denoted by $t.$
Consider a state of the clock with momentum-space wave function supported entirely in positive $\pi.$ In $\tau$-space the wave packet will move rigidly toward larger $\tau$ without changing shape. For such wave packets we can identify the clock time with Newtonian time. Such a clock is a “forward-going–clock” (FGC).
$\tau=t\ \ \ \ \ \text{FGC}$ (2.12)
Now let us combine the clock with the system and replace Newtonian time $t$ with clock time $\tau$ in the various oscillator equations. Since $\tau=t$ for our FGC clock this is trivial (just replace $t$ by $\tau$). In particular equation (2.9) becomes
$\langle 0|C|0\rangle=\frac{i}{2}\sin{\tau}$ (2.13)
Now let’s consider a time-reversed “backward-going-clock” with momentum space wave function supported in the negative $\pi$ region. For this BGC,
$\tau=-t\ \ \ \ \ \text{BGC}$ (2.14)
Equation (2.26) is replaced by,
$\langle 0|C(\tau)|0\rangle=-\frac{i}{2}\sin{\tau}$ (2.15)
Neither of these two states,
$\text{system}\otimes\text{FGC}$
and
$\text{system}\otimes\text{BGC}$
are $\cal{T}$} invariant. Only the superposition of them is $\cal{T}$}-invariant. Calling such a $\cal{T}$}invariant state $|\psi\rangle$ we may write it in the form,
$|\psi\rangle=|0\rangle\otimes\big(\ |FGC\rangle+|BGC\rangle\ \big).$ (2.16)
The state $|\psi\rangle$ is invariant under the full time-reversal in which both the system and the clock are reversed.
It is easily seen that
$\langle\psi|C(\tau)|\psi\rangle=0.$ (2.17)
In the fully time-reversed state in which both the system and the clock are reversed the expectation value of the $\cal{T}$}-odd operator in the $\cal{T}$}-even state is indeed zero.
Let me emphasize that it is not enough to time reverse the clock together with the system but we also need to replace ideal time $t$ by the clock time $\tau$ (see [1]).
2.3 Time Reversal of the Clock
One might think that under $\cal{T}$} the clock time $\tau$ should be odd but that’s incorrect. $\tau$ is the position of a particle or the hand of a clock. As such it is $\cal{T}$} invariant  [^4] In the quantum mechanics of a particle the position coordinates are invariant under time-reversal and the momentum changes sign..
${\cal{T}}\tau{\cal{T}}=\tau.$ (2.18)
The momentum $\pi$ is odd,
${\cal{T}}\pi{\cal{T}}=-\pi$ (2.19)
and so is the velocity of the clock,
$\dot{\tau}$ $=$ $-i\big[\tau,H\big]$ (2.20)
$=$ $-i\big[\tau,|\pi|\big]$ (2.22)
$=$ $\frac{\pi}{|\pi|}.$ (2.24)
2.4 Projection Operators
Before going on I want to define some notation which will come in handy when we apply these lessons to de Sitter space. Let $\Pi_{f}$ be a projection operator onto states with a FGC. In other words it projects onto states with positive values of $\pi$ while doing nothing to the system. Similarly $\Pi_{b}$ projects onto states with a BGC—negative values of $\pi$. The sum $\Pi_{f}+\Pi_{b}$ is just the unit operator. It is obviously $\cal{T}$} symmetric.
We also introduce the difference operator $\Pi_{-},$
$\Pi_{f}-\Pi_{b}=\Pi_{-}$ (2.25)
$\Pi_{-}$ is a difference of two non-overlapping projection operators. It is odd under $\cal{T}$}.
Using projection operators we can write equations (2.26) and (2.15) in the compact form,
$\langle C\rangle=\frac{i}{2}\Pi_{-}\sin{\tau}$ (2.26)
Since $\Pi_{-}$ is odd under $\cal{T}$} its expectation, as well as the expectation value of $C$ vanishes in any $\cal{T}$}-invariant state.
To summarize: There are two versions of time-reversal; one in which we reverse the system but not the clock, the second in which we reverse everything including the clock. In the first version $\langle C\rangle\neq 0,$ in the second $\langle C\rangle=0$ for any $\cal{T}$}-invariant state.
3 Spontaneous Symmetry Breaking
I will illustrate spontaneous symmetry breaking by the simple example of the 2D Ising model. Take a very large $N\times N$ lattice ($N>>1$) populated by spins shown in figure 1 for $N=15.$
Figure 1:
The lattice is the analog of the static patch of de Sitter space; the center of the lattice playing the role of the pode and the boundary playing the role of the horizon.
The Hamiltonian is
$H=\sum_{\text{nn}}Z_{i}Z_{j}+\epsilon\sum_{i}X_{i}\ \ \ \ \ (\epsilon<<1)$ (3.1)
where $\sum_{nn}$ means sum over nearest neighbors. The Hamiltonian is invariant under the discrete ${\cal{Z}}_{2}$ symmetry $K,$
$KZ_{i}K$ $=$ $-Z_{i}$ (3.2)
$KX_{i}K$ $=$ $X_{i}$ (3.4)
$KY_{i}K$ $=$ $-Y_{i}.$ (3.6)
The eigenstates of of $Z$ will be labeled by $|u\rangle$ and $|d\rangle$ (up and down). By definition,
$Z|u\rangle$ $=$ $|u\rangle$ (3.7)
$Z|d\rangle$ $=$ $-|d\rangle$ (3.9)
The ground state for $\epsilon<<1$ has the ${\cal{Z}}_{2}$ invariant form of a “cat state.”
$|gs\rangle=|uuu\cdots u\rangle+|ddd\cdots d\rangle=|U\rangle+|D\rangle$ (3.10)
3.1 Breakdown of Clustering
Consider two very distant points pandhp\ $\text${and}\ h separated by a lattice distance $\sim N.$ To be definite let $p$ be at the center and $h$ be out near the boundary as in figure 1.
Call the spin at $p$ the “observer” and the region near $h$ the laboratory, i.e., the place where experiments take place. The laboratory may well contain a genuine observer (no quotes) but that is incidental.
One finds
$\langle gs|Z_{p}|gs\rangle$ $=$ $0$ (3.11)
$\langle gs|Z_{h}|gs\rangle$ $=$ $0$ (3.13)
$\langle gs|Z_{p}Z_{h}|gs\rangle$ $=$ $1.$ (3.15)
Given that pandhp\ $\text${and}\ h are arbitrarily far apart, equations (3.15) imply a violation of cluster decomposition. States that violate cluster decomposition such as the cat state (3.10) are not normally seen in nature; they are highly unstable and tend to collapse, either to $|uuu\cdots u\rangle$ or $|ddd\cdots d\rangle.$
Restoring cluster decomposition can be done in a number of equivalent ways, all of which pick out the “observer” as a reference point to which other spins can be compared (dressed). The first way is to introduce an addition to the Hamiltonian which only involves $p.$
$H=\sum_{\text{nn}}Z_{i}Z_{j}+\epsilon\sum_{i}X_{i}+bZ_{p}$ (3.16)
In the $N\to\infty$ limit if $b$ remains finite, no matter how small, it will have a dramatic global effect on the ground state even though the perturbation only involves one spin. Depending on the sign of $b$ the new ground state will either be
$|U\rangle=|uuu\cdots u\rangle$
or
$|D\rangle=|ddd\cdots d\rangle.$
Both of these states individually satisfy cluster decomposition.
The symmetric ground state (3.10) is unstable to a tiny perturbation which not only polarizes the “observer” spin but the entire lattice. For $N\to\infty$ the magnitude of the effect becomes independent of the magnitude of $b.$
3.2 Gauging ${\cal{Z}}_{2}$
Adding an extra symmetry breaking as in (3.16) is artificial device which has no analog in de Sitter space. Another way to accomplish the same result—cluster decomposition—is to gauge gauge ${\cal{Z}}_{2}$ by restricting the physical space of states and the observables to be ${\cal{Z}}_{2}$ invariant. As before we single out $p$ as the “observer” and “dress” all observables to $p.$ This simply means replace all other  [^5] In this context the index $i$ runs over all spins other than the observer $Z_{p}.$ $Z_{i}$ by the dressed operators $\bar{Z}_{i},$
$\bar{Z}_{i}=Z_{p}Z_{i}$ (3.17)
Correlations of the gauge invariant dressed operators satisfy cluster decomposition and are identical to the correlations in the state $|U\rangle.$
3.3 Projection Operators
Let us introduce projection operators $\Pi_{u}$ and $\Pi_{d}$ which project onto up and down states of the “observer.” They satisfy the following identities,
$\Pi_{u}^{2}$ $=$ $\Pi_{d}^{2}=1$ (3.18)
$\Pi_{u}\Pi_{d}$ $=$ $0$ (3.20)
$\Pi_{u}+\Pi_{d}$ $=$ $1$ (3.22)
They also act as the identity on all spins other than $p.$
We also introduce the ${\cal{Z}}_{2}$-odd operator,
$\Pi_{-}=\Pi_{u}-\Pi_{d}$ (3.23)
Note that $\langle gs|Z_{i}|gs\rangle=0$ but $\langle gs|Z_{i}\Pi_{-}|gs\rangle\neq 0.$
4 Time-reversal in de Sitter Space
We will now apply the lessons of the previous sections to de Sitter Space, in particular to the paradox described in [7].
4.1 Static Patch Holography
[…]
$\langle A\rangle=\rm Tr\it A$ (4.1)
The degrees of freedom are in static thermal equilibrium and the only things that happen are thermal fluctuations.
The Hamiltonian is assumed to be time-reversal invariant.
Also assume an approximate dictionary relating the holographic degrees of freedom to bulk field degrees of freedom $A(r,\Omega,t).$ (Here $r,\Omega$ and $t$ are the usual static patch radial, angular and time coordinates.)
4.2 Clocks in the Static Patch
In the simple model of section 2 we had the option of defining time reversal with or without reversing the clock. In de Sitter space we don’t have that option; clocks are material systems composed of the same degrees of freedom that comprise everything else. Moreover they interact with everything else, if no other way, gravitationally. The concept of a time-reversal operation that only acts on the system, and not on the clock is at best an imprecise one. From now on $\cal{T}$} operates on all degrees of freedom including any clocks that happen to be present.
Clocks are “lumps” of matter which we assume are localized near the pode of a static patch (see however section 6). Like any other features they are rare thermal fluctuations which are emitted by the stretched horizon and are eventually reabsorbed. The probability to find a lump of mass $M$ is,
Prob $=$ $\exp{\left(-\frac{M}{T_{gh}}\right)}$ (4.2)
$=$ $\exp{\left(-2\pi\ell M\right)}$ (4.4)
where $\ell$ is the de Sitter radius and the subscript $gh$ refers to Gibbons-Hawking.
It is evident that lumps which can reasonably called clocks are extremely (exponentially) rare, but one can condition of the existence of a clock. A formal way to do that is to introduce projection operators in the static patch Hilbert space. The projection operators are bulk operators that in principle can be expressed in terms of the holographic degrees of freedom. The details are not important other than from the bulk viewpoint they are localized near the pode, far from the horizon.
Let $\Pi$ be a $\cal{T}$}-symmetric projection operator that projects onto a subspace in which a clock of some particular type is present at the pode. I will assume that the clock can be either forward or backward going, and that $\Pi$ is symmetrically composed of two non-overlapping projection operators: a forward-going and a backward going projection. We also define $\Pi_{-},$
$\Pi$ $=$ $\Pi_{+}=\Pi_{f}+\Pi_{b}$ (4.5)
$\Pi_{f}$ $=$ ${\cal{T}}\Pi_{b}{\cal{T}}$ (4.7)
$\Pi_{-}$ $=$ $\Pi_{f}-\Pi_{b}$ (4.9)
Note that $\Pi_{-}$ is not a projection operator; its square satisfies,
$\Pi_{-}^{2}=\Pi\neq 1.$ (4.12)
As in section 2 we will use the notation $\tau$ for clock time. For a FGC
$\tau=t$
and for a BGC
$\tau=-t.$
Conditioning on the existence of a clock without regard to whether it is a FGC or BGC is accomplished by replacing the maximally mixed density matrix by $\Pi$. Expectation values are given by,
$\langle A\rangle=\rm Tr\it\Pi A=\rm Tr\it\big(\Pi_{f}+\Pi_{b}\big)A$ (4.13)
We may also condition on a FGC (BGC) by applying $\Pi_{f}$ ($\Pi_{b}$).
$\langle A\rangle$ $=$ $\rm Tr\it\Pi_{f}A$ (4.14)
$=$ $\frac{1}{2}\rm Tr\it\ \big(\Pi_{-}A\big)$ (4.16)
So far this parallels section 2.2. But as a bonus it also parallels section 3.1; the location of the operators $\Pi,\Pi_{-},\Pi_{f},\Pi_{b}$ is at the pode—the center of the static patch analogous to the center of the square lattice in figure 1. The operator $A$ is out near the horizon analogous to the point $h.$
In parallel with (4.23) we also consider the operator
C=i[A(s),A(s)]C=i[A(-s),A(s)]
for a fixed numerical value of a time $s.$ $C$ is odd under $\cal{T}$}.
The expectation value of $C$ is an order parameter for SSB of time-reversal analogous to $Z_{i}$ being an order parameter for the ${\cal{Z}}_{2}$ symmetry $K$ in the Ising model.
Since $\Pi_{-}$ and $C$ are $\cal{T}$}-odd their expectation values vanish in the maximally mixed state,
$\rm Tr\it\Pi_{-}$ $=$ $0$ (4.17)
$\rm Tr\it C$ $=$ $0$ (4.19)
The distance between the pode and the horizon becomes infinite in the limit of infinite de Sitter radius, but according to assumption the correlation between $\cal{T}$} and $\Pi_{-}$ does not go to zero; $\langle C\ \Pi_{-}\rangle$ tends to a finite limit (see the introduction section 1). Therefore cluster decomposition is violated in the bulk theory, signaling a spontaneous breaking of $\cal{T}$}.
Simply stated, the existence of a tiny time-oriented feature, arbitrarily far from the laboratory near the horizon will set off an instability that leads to a non-zero expectation value for $C,$ consistent with the non-zero value in the semiclassical theory, and cluster decomposition is restored [^6] While this eliminates the paradox it still remains to calculate the values of the correlations in a specific holographic theory such as $\rm DSSYK_{\infty}$} and show that they agree with the semiclassical theory..
One feature of SSB is that the instability leading to a properly clustering state can be kicked off by a very weak perturbation. In fact almost any perturbation will work and lead to the same global behavior. This meshes well with the literature in which the “observer” can be almost anything from a galaxy to a few low energy photons.
4.3 The Puzzle
Consider a bulk operator $A(t)$ which for definiteness is located in the laboratory: the region near, but not too near the horizon. $A$ is Hermitian and time-reversal symmetric,
$A$ $=$ $A^{{\dagger}}$ (4.20)
${\cal{T}}A(t){\cal{T}}$ $=$ $A(-t).$ (4.22)
The operator
C=i[A(s),A(s)]C=i[A(-s),A(s)] (4.23)
is $\cal{T}$} odd for any value of $s.$
The expectation value of $C$ vanishes in the maximally mixed state,
$\rm Tr\it C=i\rm Tr\it[A(-s),A(s)]=0.$ (4.24)
The reason is just that the trace of a commutator vanishes.
Equation (4.24) has presented a puzzle for some time [7], the reason being that it is not zero in the semiclassical theory of de Sitter space. The quantity $\langle C\rangle$ is equal to the imaginary part of the correlation function $\langle A(-t)A(t)\rangle.$ When calculated in QFT in a de Sitter background the imaginary part of two-point functions is not zero [^7] At this point the reader may want to review sections 2.1 and 2.2.
This is one of those puzzles whose resolution requires the existence of a time-oriented clock—let’s say a FGC. While (4.24) is correct the presence of a FGC breaks $\cal{T}$} and allows a non-vanishing value of $\langle C\rangle.$ To see this, observe that both $C$ and $\Pi_{-}$ are odd under time reversal so that $C\Pi_{-}$ is even. This allows
$\langle C\Pi_{-}\rangle=\rm Tr\it C\Pi_{-}\neq 0.$ (4.25)
These observations: the need to measure time by a physical FGC, but that is not enough. In order that the clock can be arbitrarily far from the location of $C$ and (4.25) remain finite there must be a spontaneous breaking of $\cal{T}$} in the maximally mixed state.
Physical clocks and spontaneous breaking of $\cal{T}$} can potentially resolve the paradox of [7]. Simply stated, the existence of a tiny time-oriented feature, arbitrarily far from the laboratory (near the horizon) will set off an instability that leads to a non-zero expectation value for $C,$ consistent with the non-zero value in the semiclassical theory.
5 Gauging $CT$ and Dressing C
In section 3.2 I discussed gauging ${\cal{Z}}_{2}$ as a way to restore cluster decomposition. In de Sitter space this may be much more natural than adding a small symmetry breaking to the holographic Hamiltonian. Harlow and Numasawa [6] have made arguments that $\cal{T}$} is a gauge symmetry which I find compelling. In the ${\cal{Z}}_{2}$ case we constructed gauge invariant operators by dressing to the “observer,”
$\bar{Z}_{i}=Z_{i}Z_{p},$
which rendered them gauge invariant and properly clustering. While $Z_{i}$ is odd under $K$ and is therefore not gauge invariant, $\bar{Z}_{i}$ is invariant.
The analog of $Z_{i}$ is the order parameter $C.$ Like $Z_{i}$ it is odd (under $\cal{T}$}) but it can be dressed to the clock:
$\bar{C}=C\Pi_{-}$ (5.1)
$\bar{C}$ is the dressed version of $C;$ it is $\cal{T}$}-invariant. Unlike correlations of $C,$ correlations of $\bar{C}$ do cluster properly.
If, following [6] we take $\cal{T}$} to be a gauge symmetry then $C$ is not gauge invariant, but the dressed operator $\bar{C}$ is.
6 Another Viewpoint
It is interesting to consider the flat-space limit in which $\ell\to\infty$ and the pode moves infinitely far from the horizon. One can take the flat-space limit in two ways. One way is to focus attention on the near horizon region which includes everything whose proper distance from the horizon remains finite in the limit $\ell\to\infty.$ This is the limit in which the static patch tends to Rindler space and the pode disappears to infinity.
The other way to take the limit is by focusing attention on the pode while the horizon moves off to infinity. The pode just becomes the origin of Minkowski space.
So far I have thought of the “observer” as living at the pode and the laboratory including $C$ as being near (but not too near) the horizon. This seems strange because one would expect that when the observer moves off to infinity its presence should become irrelevant to the physics in the laboratory. The resolution of this puzzle—SSB—is what this paper is all about. But it is also interesting to center the the laboratory at the pode and let the “observer” follow the horizon (or stretched horizon).
In this latter viewpoint the clock is a tiny out-of-equilibrium feature near the horizon (in the Hawking-Unruh radiation) as long as it breaks time-reversal symmetry. (Breaking $\cal{T}$} is very generic; not breaking it requires fine-tuning.)
Consistency requires the following conjecture to be true:
Conditioning on such a small horizon asymmetry correlation functions near the pode, calculated in the holographic theory, agree with their semi-classical counterparts.
I emphasize that this is a conjecture; I don’t see any obstacle to it being true but it needs confirmation.
7 Conclusion
Depending on one’s definition of observer, the claim that quantum de Sitter space is consistent only if there is an observer is either wrong, or the definition of observer is so broad that it means almost anything. What is true is that the $\cal{T}$} symmetric vacuum is unstable and only the feeblest perturbation will suffice to kick it into a stable spontaneously broken state. That perturbation could be the existence of a particle detector, a gravitationally bound galaxy, a tiny black hole or a couple of photons of energy $T_{gh}$ as long as it is not exactly $\cal{T}$}-symmetric. For all but the last case the probability for the existence of such an object is exponentially small,
$\sim\exp{\left(-\frac{M}{T_{gh}}\right)}=\exp{\left(-2\pi\ell M\right)}$
but for the last case where the energy of the feature is $\sim T_{gh}$ the probability is order unity. One can summarize this with the phrase, “Observers” are common; Observers are rare.
In fact it is probably true that the theory without an “observer” is inconsistent; a state with a feeble observer is not orthogonal to the maximally mixed state. Projecting onto a state in which every conceivable “observer” is absent probably leaves nothing. But states without observers are consistent. If for example an observer means anything with mass greater than the Planck scale the probability to not have an observer is close to one,
$\text{Prob for no observer}=1-\left(e^{-\frac{M_{planck}}{T_{gh}}}\right)$
By contrast, the probability to not have an “observer” of any kind may be zero.
Finally we can ask about stability of the vacuum: does it require “observers” or to put it another way is the Hartle-Hawking vacuum unstable. Its holographic dual—the maximally mixed state is unstable. The answer, based on Euclidean gravitational path integration is yes, the HH state is unstable. The argument was originally due to Polchinski [8]. Recently it was revisited in [3] and [4]. The conclusion of these later papers is that the presence of an observer in the path integral is sufficient to stabilize the vacuum. It is almost certainly true that a feeble “observer” will also suffice and I will assume that this is the case.
Whether the Polchinski instability is related to the SSB instability discussed in this paper is an open question. Both instabilities are eliminated by the introduction of an “observer.”
I want to emphasize a point made in [1]. It is not enough to project onto states with an clock—FGC or BGC—to reproduce semiclassical correlations. It is also necessary to replace the ideal time $t$ by the clock time $\tau$ as the argument of the correlation functions.
Finally a question: If $\cal{T}$} is a spontaneously broken discrete gauge symmetry (Higgs phenomenon) are there consequences of this? In particular does parallel transport around the bifurcate horizon induce a $\cal{T}$} reversal?
Acknowledgement
I have benefited from many discussions with Ying Zhao, Daniel Harlow and Edward Witten.